# Integrating anatomy and electrophysiology in the healthy human heart: Insights from biventricular statistical shape analysis using universal coordinates

### v1

@@ 3.6 Statistical analyses
The correlation analysis between PC scores and ECG-derived features revealed eight statistically significant correlations between biventricular anatomy and EP in our study population, of which the strongest and most significant one was the correlation between PC1 and RRint, with a correlation coefficient equal to -0.50. Furthermore, six statistically significant correlations were found between anatomical components and numerical demographic features, the strongest ones being the correlation between PC1 and height, followed by PC1 and BSA, with correlation coefficients equal to -0.62 and -0.55, respectively. The complete list of correlation coefficients is available in Table 3. The difference between the scores of PC1 was significant between all demographic groups that were compared in statistical tests, except between the male early athletes and the male late athletes. For PC2 and PC3, no such statistically significant differences between demographic groups was identified. The relevant distributions are shown in A.1.
@@ 3.7 Available material
The biventricular SSM, including PCs, derived from the full population, will be made publicly available on Zenodo doi:10.5281/zenodo.14261122, along with a synthetic cohort of 100 biventricular, volumetric meshes, including fiber orientation and anatomical tags for LV and RV endo- and epicardium, and LV septum. The volumetric meshes of the synthetic cohort will be provided in VTK format. Code to generate surface meshes from the published SSMs, by sampling PC scores, will be made available in the Github repository https://github.com/LoreVanSantvliet/BiventricularSSM.
@@ 4.3 Biventricular anatomy in demographic groups
Statistically significant differences were also observed in PC1 between the healthy control group, and both early and late athletes, indicating that both groups of male athletes have significantly larger ventricles compared to male subjects in the healthy control group. This finding is consistent with existing literature on athletic hearts. Morganroth et al. reported an increased LV end-diastolic volume and mass for athletes engaging in isotonic exercise, compared to healthy controls [57]. Similarly, Oxborough et al. found that RV chamber dimensions in athletes often exceed normal ranges, particularly towards the large end of the spectrum [58].
@@ 4.4 Sex-related anatomical differences
There is a growing interest in sex-related differences in cardiac anatomy and function, as these differences are crucial for enhancing cardiovascular care and ensuring equitable treatment across sexes [59, 60, 45, 61]. We performed a range of experiments to assess the impact of sex on the performance and representativeness of the SSM. As described in Section 4.3, one key finding was the significant difference in biventricular size that we observed between female subjects of our study population, and male subjects belonging to the healthy control group.
@@ 4.5 Impact of dataset size
We investigated the reconstruction error in a series of LOOCV experiments varying the SSM training set size. The decreasing trend in reconstruction error when increasing the training set size is in line with the findings from Bai et al. [13], who showed a substantial reduction in reconstruction error for ventricular geometries with dataset sizes up to 400-600 subjects, when only the first 100 PCs were considered for reconstruction. Our findings, combined with those of Bai et al., suggest that a straightforward improvement to our model would be to include an even larger number of healthy subjects. This could not only further reduce reconstruction error, but also increase the statistical power when correlating PC scores to other health-related variables, possibly enabling the identification of additional, significant sex-related differences between PCs.
@@ 4.6 Correlations
The positive correlation between heart size and RR interval is consistent with prior research: LV mass index has been related to the RR interval through correlation with the heart rate [65]. In electromechanical simulations, Rodero et al. [15] observed a positive correlation between the second mode of variation in their SSM, encoding heart size, and the duration of systole, closely related to the QT interval. This supports our findings that larger hearts tend to have longer QT intervals. We did not, however, find studies validating the correlation between PC1 and T wave duration.
@@
The previously described correlations of height, weight and BSA with PC1 aligns with the work of Hense et al. [66], who identified significant positive correlations between LV mass on the one hand, and weight, height, and BMI on the other hand. However, LV mass and BSA correlation were not investigated. Second, in a series of multivariate linear regression analyses, Pfaffenberger et al. [56] examined the influence of demographic factors on a range of variables related to ventricular size (LV end-diastolic and end-systolic diameter, LV end-diastolic and end-systolic volume, RV end-diastolic diameter, RV end-diastolic area). They found significant effects of height, weight, and BSA on most of these variables, consistently with our results.
4.7 CDT validation
Validation of CDT technology is highly challenging [5]. How can we confirm that a set of PDEs adequately reflects the physiological behavior of the cardiovascular (sub)system under investigation for a given application? Validation approaches for cardiac digital twins in healthcare rely on different types of evidence [67]. The first category consists of evidence related to the mathematical equations employed, including modeling assumptions and parameters. The second category, calibration evidence, involves demonstrating alignment between the model’s simulated or predicted results and the clinical data used for calibration. Finally, validation evidence pertains to the model’s ability to accurately replicate real-world data that was not used at any stage of model construction.
@@ 5 Conclusion
This study reports on a highly automated workflow for the construction of SSM of biventricular anatomy. The training set was built upon high-resolution cardiac CT scans of 271 healthy subjects, including athletes. A novel, lightweight UVC-based approach for establishing correspondence between geometries was introduced. The resulting SSM was used in various downstream analyses revealing quantitative relationships between ventricular anatomy and ECG-derived features, as well as demographic features. Additionally, a virtual cohort of 100 volumetric, biventricular meshes, designed for use in electrophysiological simulations, was extracted from the SSM and published.
@@
This research is funded by a PhD fellowship fundamental research from the Research Foundation - Flanders, Belgium (FWO) for L. Van Santvliet (FWO project number 1107725N); the Flanders AI Research Program, Belgium; HORIZON-HLTH-2022-IND-13: ‘Privacy compliant health data as a service for AI development (PHASE IV AI)’, European Union (grant agreement no. 101095384); a research grant from the Research Foundation - Flanders, Belgium (FWO) for R. Willems (FWO project number T003717N). This research received support from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant TwinCare-AF (grant agreement no. 101148636) for E. Zappon, who also acknowledges her membership to INdAM GNCS - Gruppo Nazionale per il Calcolo Scientifico (National Group for Scientific Computing, Italy). Further financial support was received from the Austrian Research Promotion Agency (FFG) within the InstaTwin project, grant FO999891133.
Appendix
A.1 Impact of sex-related anatomical differences
@@
A.2 Construction of synthetic cohort

### v2

@@ 3.6 Statistical analyses
The correlation analysis between PC scores and ECG-derived features revealed eight statistically significant correlations between biventricular anatomy and EP in our study population, of which the strongest and most significant one was the correlation between PC1 and RRint, with a correlation coefficient equal to -0.50. Furthermore, six statistically significant correlations were found between anatomical components and numerical demographic features, the strongest ones being the correlation between PC1 and height, followed by PC1 and BSA, with correlation coefficients equal to -0.62 and -0.55, respectively. The complete list of correlation coefficients is available in Table 3. The difference between the scores of PC1 was significant between all demographic groups that were compared in statistical tests, except between the male early-onset athletes and the male late-onset athletes. For PC2 and PC3, no such statistically significant differences between demographic groups was identified. The relevant distributions are shown in D.
@@ 3.7 Available material
The biventricular SSM, including PCs, derived from the full population, is made publicly available on Zenodo doi:10.5281/zenodo.14261122, along with a synthetic cohort of 100 biventricular, volumetric meshes, including fiber orientation and anatomical tags for LV and RV endo- and epicardium, and LV septum. The volumetric meshes of the synthetic cohort are provided in VTK format. Code to generate surface meshes from the published SSMs, by sampling PC scores, is made available in the Github repository https://github.com/LoreVanSantvliet/BiventricularSSM.
@@ 4.3 Biventricular anatomy in demographic groups
Statistically significant differences were also observed in PC1 between the healthy control group, and both early- and late-onset athletes, indicating that both groups of male athletes have significantly larger ventricles compared to male subjects in the healthy control group. This finding is consistent with existing literature on athletic hearts. Morganroth et al. reported an increased LV end-diastolic volume and mass for athletes engaging in isotonic exercise, compared to healthy controls [61]. Similarly, Oxborough et al. found that RV chamber dimensions in athletes often exceed normal ranges, particularly towards the large end of the spectrum [62].
@@ 4.4 Sex-related anatomical differences
There is a growing interest in sex-related differences in cardiac anatomy and function, as these differences are crucial for enhancing cardiovascular care and ensuring equitable treatment across sexes [63, 64, 49, 65]. We performed a variety of experiments to assess the impact of sex on the performance and representativeness of the SSM. As described in Section 4.3, one key finding was the significant difference in biventricular size that we observed between female subjects of our study population, and male subjects belonging to the healthy control group. This sex-related difference is consistent with findings by Burns et al. [20].
@@ 4.5 Impact of dataset size
We investigated the reconstruction error in a series of LOOCV experiments varying the SSM training set size. The decreasing trend in reconstruction error when increasing the training set size is in line with the findings of Bai et al. [16], who showed a substantial reduction in reconstruction error for ventricular geometries with dataset sizes up to 400-600 subjects, when only the first 100 PCs were considered for reconstruction. Our findings, combined with those of Bai et al., suggest that a straightforward improvement to our model would be to include an even larger number of healthy subjects. This could not only further reduce reconstruction error, but also increase the statistical power when correlating PC scores to other health-related variables, possibly enabling the identification of additional, significant sex-related differences between PCs.
4.6 Synthetic cohort and CDTs
Leveraging the diversity captured by our biventricular SSM, we created a synthetic cohort of anatomically detailed biventricular meshes. The geometries in this cohort were each annotated with anatomical labels and fiber orientations. This cohort, and possibly even more synthetic geometries derived from our SSM, can be used for large-scale simulation studies. In addition to the fundamental insights that large-scale simulation studies can provide on simulation technology, for example by clarifying their parameter sensitivities and uncertainties [66], they can also be used to create simulation databases [23]. Simulation databases, in turn, are a useful resource for training data-driven calibration models to perform automated functional personalization (calibration) of patient-level CDTs. A simulation database suitable for training advanced, anatomy-aware, data-driven calibration models, such as DL models, needs to be large and diverse. Both of these requirements are made more accessible by the release of our SSM, which offers anatomical diversity without requiring researchers to access or process sensitive health data, and which offers the ability to generate a virtually unlimited amount of synthetic, yet realistic heart shapes.
Additionally, the synthetic cohort itself serves as an important building block toward population-level CDTs. Such CDTs, also called virtual cohorts, have application potential in virtual clinical trials [12, 13]. We envision the use of our anatomical cohort of ready-to-use biventricular meshes, representative of a healthy population, to create a healthy virtual control population for such trials.
@@ 4.7 Correlations
The positive correlation between heart size and RR interval is consistent with prior research: LV mass index has been related to the RR interval through correlation with the heart rate [69]. Additionally, Burns et al. also detected a significant negative correlation between heart size and heart rate [20]. In electromechanical simulations, Rodero et al. [18] observed a positive correlation between the second mode of variation in their SSM, encoding heart size, and the duration of systole, closely related to the QT interval. This supports our findings that larger hearts tend to have longer QT intervals. We did not, however, find studies validating the correlation between PC1 and T wave duration.
@@
The significant correlation between a principal component encoding ventricular size and height is confirmed in [20]. Both the correlation between ventricle elongation and BMI, and between heart size and age are consistent with findings of Burns et al. [20] as well. Moreover, the correlations of height, weight and BSA with PC1 aligns with the work of Hense et al. [70], who identified significant positive correlations between LV mass on the one hand, and weight, height, and BMI on the other hand. However, LV mass and BSA correlation were not investigated in this study. Finally, in a series of multivariate linear regression analyses, Pfaffenberger et al. [60] examined the influence of demographic factors on a range of variables related to ventricular size (LV end-diastolic and end-systolic diameter, LV end-diastolic and end-systolic volume, RV end-diastolic diameter, RV end-diastolic area). They found significant effects of height, weight, and BSA on most of these variables, consistent with our results.
4.8 Applications in patient-level CDTs
Our automated workflow for anatomical twinning is directly applicable for generating patient-level CDTs. Beyond this methodological applicability, the insights that are drawn from our statistical analysis of the relationship between ventricular anatomy and electrophysiology, are more broadly useful for CDT research. As an example, we propose their use for establishing which anatomical factors are most important to capture in a CDT. For example, statistical correlations revealed in this work, presented in Table 3, confirm that especially heart size is very important to model accurately in CDTs, as it has a significant effect on many different aspects of clinical ECGs.
A second possible application of the statistical insights is in validation of CDT simulation technology. Validation of CDT technology is highly challenging [5]. How can we confirm that a set of PDEs adequately reflects the physiological behavior of the cardiovascular (sub)system under investigation for a given application? Validation approaches for cardiac digital twins in healthcare rely on different types of evidence [71]. The first category consists of evidence related to the mathematical equations employed, including modeling assumptions and parameters. The second category, calibration evidence, involves demonstrating alignment between the model’s simulated or predicted results and the clinical data used for calibration. Finally, validation evidence pertains to the model’s ability to accurately replicate real-world data that was not used at any stage of model construction.
@@ 4.8 Applications in patient-level CDTs
4.9 Limitations and future extensions
For male subjects, data was derived from the Master@Heart study, a prospective cohort trial including athletes and healthy control subjects [32]. For female subjects, on the other hand, data was obtained from a clinical referral of the subjects to a cardiac CT. These subjects can still be considered healthy, since they are not diagnosed with any cardiovascular disease, and subjects that presented with more than three cardiovascular risk factors were excluded. However, their healthy status might still not match that of subjects in the Master@Heart study, with more selective exclusion criteria.
Our SSM covers biventricular cardiac anatomy below the cardiac valves, excluding the non-conductive cardiac skeleton. Given the electrophysiological focus of our work, this non-conductive tissue is less important to incorporate into our SSM. However, for enabling the use of the provided meshes for mechanical or hemodynamical modeling, valvular anatomy should be included. In this case, an adapted UVC system that includes cardiac valves, such as the one provided in [72], could be valuable for extending our UVC-based approach to establish correspondence between geometries to valvular structures. If the extended UVCs are ordered in a sufficiently regular fashion in Cartesian space to enable subsequent triangulation, as covered in section 2.2.1, an extension of the SSM towards additional anatomical structures would then be straightforward. However, incorporating additional structures, and hence possibly introducing additional anatomical variability is expected to further increase data requirements for obtaining a representative SSM.
Additionally to the ventricles, our SSM could be extended towards the atrial structures as well, leading to a four-chamber SSM. Given the importance of the atria in establishing P wave morphology, this would be an interesting extension, allowing further investigation of relationships between cardiac anatomy and electrophysiology. Given that atria are structurally highly complex [19], arguably even more so that ventricles, it is expected that a larger dataset would be more suitable for this extension. To include the atria, the correspondence establishment could be extended by making use of universal atrial coordinates [30, 31] in addition to UVCs. Similarly to before, the ordering of these coordinates needs to be sufficiently regular as to allow subsequent triangulation, possibly with the aid of Laplacian smoothing. A suitable sampling scheme over the different atrial coordinates needs to be devised, leading to an additional number of points per mesh, thus an additional number of input features for the PCA.
The results described in Section 3.3 reveal that increasing the dataset would likely lead to further improvements of the SSM. The UK Biobank, used in [17, 20], is a suitable resource for increasing the dataset size, although the imaging resolution of these cardiac MR images is lower than the resolution of CT scans used in this study.
@@ 5 Conclusion
This study reports on a highly automated workflow for the construction of an SSM of biventricular anatomy. The training set was built upon high-resolution cardiac CT scans of 271 healthy subjects, including athletes. A novel, lightweight UVC-based approach for establishing correspondence between geometries was introduced. The resulting SSM was used in various downstream analyses revealing quantitative relationships between ventricular anatomy and ECG-derived features, as well as demographic features. Additionally, a virtual cohort of 100 volumetric, biventricular meshes, designed for use in electrophysiological simulations, was extracted from the SSM and published.
@@
This research is funded by a PhD fellowship fundamental research from the Research Foundation - Flanders, Belgium (FWO) for L. Van Santvliet (FWO project number 1107725N); the Flanders AI Research Program, Belgium; HORIZON-HLTH-2022-IND-13: ‘Privacy compliant health data as a service for AI development (PHASE IV AI)’, European Union (grant agreement no. 101095384); a research grant from the Research Foundation - Flanders, Belgium (FWO) for M. De Vos (FWO project number G046925N); a research grant from the Research Foundation - Flanders, Belgium (FWO) for R. Willems (FWO project number T003717N). This research received support from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant TwinCare-AF (grant agreement no. 101148636) for E. Zappon, who also acknowledges her membership to INdAM GNCS - Gruppo Nazionale per il Calcolo Scientifico (National Group for Scientific Computing, Italy). Further financial support was received from the Austrian Research Promotion Agency (FFG) within the InstaTwin project, grant FO999891133, and from the Austrian Science Fund (FWF), grant 10.55776/PAT174842.
Appendix A Evaluation of UVC-based sampling
We studied the impact of the number of mesh points, termed the mesh size, on the sampling error. Starting from the original biventricular meshes, constructed as explained in Section 2.1.3, we resampled all 271 meshes of the complete dataset using a variety of sampling schemes, displayed in Table 4. By using a constant scaling factor for adapting each of the three meshing parameters (the number of values, the number of values for the LV and the number of values for the RV), the element shapes are kept constant over all meshing schemes. An identical apical coarsening procedure to the one described in Section 2.2.1 was followed for all sampling schemes. Sampling errors were defined as point-to-surface distances of resampled mesh points to the original mesh, averaged over all mesh points. They were calculated using the vtkDistancePolyDataFilter (Kitware, Clifton Park, New York, USA).
| # values | # values | # values | Size | |
| 1 | ||||
| 2 | ||||
| 3 | ||||
| 4 | ||||
| 5 |
Figure 14 shows the distribution of the sampling errors of all meshes of the complete dataset for each of the five sampling schemes. A clear trend is visible, with lower sampling errors for increasing mesh sizes. This trend is expected: the more mesh points are included in a mesh, the better we expect it to represent shapes in their full complexity. However, it is also expected that at some point, the advantage of having a larger mesh size is not as pronounced anymore, i.e. the sampling error stagnates. This justifies our choice of sampling scheme 4.
Appendix B Biventricular anatomy in balanced demographic groups
The analysis reported in Section 4.3 was repeated using a balanced dataset containing 56 subjects from each of the four demographic groups instead of the slightly imbalanced complete dataset. The results of this analysis, displayed in Figure 15 show an overall slightly increased reconstruction error due to the smaller overall dataset size: 0.59 for the female group, and 0.69 , 0.73 and 0.68 for the male control group, male early-onset athletes, and male late-onset athletes, respectively. Qualitatively, the differences in reconstruction error between the various demographic groups remain unchanged.
Appendix C Influence of shape factors independent of heart size
To reveal the effect of biventricular shape differences between demographic groups and between sexes, independent of heart size, additional analyses were performed. For these analyses, scaling was allowed in the Procrustes alignment, leading to SSMs where heart size was not encoded as one of the PCs. Apart from this change, an identical methodology to the one described in B and Section 2.2.3 was used for the demographic groups and sexes, respectively.
Figure 16 shows the reconstruction errors per demographic group. The female group was associated with the smallest reconstruction errors, although the difference was much smaller compared to Figure 15. Therefore, we conclude that the smaller reconstruction error for this demographic group can be largely but not entirely explained by heart size.
Figure 17 shows the effect of sex-related anatomical differences, excluding heart size, on representativeness of biventricular SSMs. Similarly to Figure 11, the reconstruction of female hearts was associated with lower reconstruction errors than the reconstruction of male hearts, but the difference is lower when excluding heart size from the analysis. The interpretation of representativeness of opposite-, same- or mixed-sex SSMs remains the same as that of the SSMs including heart size, laid out in Section 3.4.
Appendix D Impact of sex-related anatomical differences
@@
Appendix E Construction of synthetic cohort